OpenStudy (anonymous):

a printer cost 150 and this weekend its 30% off

OpenStudy (saifoo.khan):

First find 30% of 150.

OpenStudy (anonymous):

10%=15, 20%= 30, 30%=45. 150-45=$105

OpenStudy (saifoo.khan):

30% = 30/100 = 0.3 Now 0.3 * 150/ This will give u the discount price.

OpenStudy (anonymous):

thanks